BEN IS MAKING WOODEN FOR THE NEXT ARTS AND CRAFTS SALE. EACH TOY COSTS BEN $1.80 TO MAKE. IF HE SELLS THE TOYS FOR $3.00 EACH, HOW MANY WILL HE HAVE TO SELL TO MAKE A PROFIT OF EXACTLY $36.00?
Answer by stanbon(75887)

BEN IS MAKING WOODEN FOR THE NEXT ARTS AND CRAFTS SALE. EACH TOY COSTS BEN $1.80 TO MAKE. IF HE SELLS THE TOYS FOR $3.00 EACH, HOW MANY WILL HE HAVE TO SELL TO MAKE A PROFIT OF EXACTLY $36.00?
----------
Profit on one toy = 3.00-1.80 = 1.20
---------------
36.00/1.20 = 30 toys
He will have to see 30 toys.
=========
Cheers,
Stan H.